What number must be subtracted from -5 for the difference to be -2
1
Expert's answer
2019-10-21T10:29:54-0400

Let "X" be an unknown number.

The equation will be the next

"-5 - X = -2"

After easy arithmetic operations we will have

"-5 - (-2) = -(-X)" or

"-5+2 = X" or

"-3 = X"

Thus, the answer is "X = -3" .